3. An L 5x 3 %x % angle is used as a compression member of a bridge. If L-10 feet, determine the allowable stress for Fy -36 ksi. What boundary conditions do you pick for the two ends? Why? What is the allowable axial load for extension? Explain.

PLEASE, I WANT THE ANSWER WITH GOOD HANDWRITING PLEASE, AND CLEAN AND CLEAR. THANK YOU​
Answer #1

minimum radius of gyration of L5x3-1/2x3/4 = 0.744 in

length of angle section = 10 ft

assuming pinned condition, effective length = 10*12=120 in

slenderness ratio = 120/0.744 = 161.3

Euler buckling stress = pi2E/(KL/r)2 = pi2*29000/(161.32)=11 ksi

critical stress=0.658(36/11)*36=9.1 ksi

area of cross section=5.85 in2

allowable load on member = 5.85*9.1=53.2 kips